A quality-conscious disk manufacturer wishes to know the fraction of disks his company makes which are defective. Step 1 of 2: Suppose a sample of 985 floppy disks is drawn. Of these disks, 917 were not defective. Using the data, estimate the proportion of disks which are defective. Enter your answer as a fraction or a decimal number rounded to three decimal places.

Suppose a sample of 985 floppy disks is drawn. Of these disks, 917 were not defective. Using the data, construct the 85% confidence interval for the population proportion of disks which are defective. Round your answers to three decimal places.
Answer #1

Proportion of disks which are defective = Number of disks which are defective / Total number of disks

Total number of disks = 985

Number of disks which are defective = 985 - 917 = 68

Proportion of disks which are defective = p = 68 / 985 = 0.069

Now, we find the confidence interval for the population proportion of disks which are defective using the formula :

p(1 - p) Upper limit: p + Za/21

Lower limit: p -Z/1/p1-p)

We have p = 0.069

n = 985

a1-0.85-0.15

Using Standard Normal Tables we get :

alpha/2-Zo.15/2 Z0.075 1.44

Putting all the values in the above expression we get ,

Upper limit 0.069 1.1 0009-0.069) er (am 0.081 985

/ V 0.069(1 -0.069) 985 Lower limit = 0.069-1.44 = 0.057

So,  85% confidence interval for the population proportion of disks which are defective is : ( 0.057 , 0.081 )